what mass of calcium will produce 10g of calcium oxide


1
Expert's answer
2021-12-13T17:04:03-0500

2 Ca + O 2 = 2CaO

Moles of CaO=mass/RFM

=10/56.0774=0.17857 moles

Ca: CaO

2:2

Moles of Ca that reacted=0.17857*40

=7.143g